OpenStudy (anonymous):

Rationalize the denominator of square root of negative 9 over open parentheses 4 minus 7 I close parentheses minus open parentheses 6 minus 6 I close parentheses

OpenStudy (anonymous):

\[\sqrt{-9} / (4-7i)-(6-6i)\]

OpenStudy (ladiesman217):

solve the first parentheses and get the square root of -p which should give you 3/3=1 so its 1- (6-6i) which is 0 so its 1-0 = -1

OpenStudy (ladiesman217):

is that helpful